1.     Suppose the IS curve is Y = 39XX-100i and Y = 1500 + 250i is the LM curve, where XX is the last two digits of your ID number. Using these compute:

a)     The equilibrium interest rate and output (i*and Y*).

b)     If government spending was increased by 100m with an immediate impact elasticity of 2.5 in the goods market, determine new income and interest rate.

c)     Determine the impact of the above policy on private investment if it is known that di/dA = XX/100, where XX is the last two digits of your ID number.

d)     Determine the magnitude of the change in money supply required to eliminate any crowding out effect in (c) above. Suppose di/dMs = -0.1X, where X is the last digit of your ID number.

e)     Explain the dynamics represented in (a-d) using an IS-LM space. (You may insert a snapshot of the graph if drawn manually).


1
Expert's answer
2021-05-08T14:41:11-0400

let XX=10

a)IS=LM

"3910-100i=1500 + 250i"

i=6.89

"Y=1500 + 250i =1500+250\times6.89=3 222"

b)

"mg=\frac{\Delta Y}{\Delta G}"


"2.5=\\frac{\\Delta Y}{100}\n\n\u200b"

"\Delta Y=250"

Y=3222+250=3472

"Y=1500 + 250i"

"3472=1500 + 250i"

i=7.89

c)

"\frac{di}{dA} =\frac{10}{100}"


"\frac{7.89-6.89}{dA} =\frac{10}{100}"

dA=10

d)

"\frac{di}{dMs}= -0.10"


"\frac{7.89-6.89}{dMs}= -0.10"


dMs=-10


e)




the increase and decrease in interest rates, respectively, contributes to the reduction and growth of investments